1.   The weight of Group 1:Jay - 43 kg, Stella - 42 kg, Tilly - 44 kg, Joanna - 53 kg. Consider samples of size 3 that can be drawn from this population.

a.   How many possible samples can be drawn?

b.   List all the possible samples and the corresponding means.  

c.    Construct the sampling distribution of the sample means. 



1
Expert's answer
2022-03-11T02:12:43-0500

Given N "=4" and sample size "n=3"


(a) Possible samples

"=(_n^N)=(_3^4)=4" samples


(b) Possible samples and their means


"\def\arraystretch{1.5}\begin{array}{c:c:c} no.& Sample &Mean\\\hline1.&42,43,44&43\\\hdashline2.&42,43,53&46\\\hdashline3.&42,44,53&46.333\\\hdashline4.&43,44,53&46.667\\\hline\end{array}"


(c) Sampling distribution of the sample means

"\def\arraystretch{1.5}\begin{array}{c:c:c:c} \bar{X}&f&f(\bar{X})&\bar{X}f(\bar{X})\\\hline43&1&\frac{1}{4}&\frac{43}{4}\\\hdashline46&1&\frac{1}{4}&\frac{46}{4}\\\hdashline\frac{139}{3}&1&\frac{1}{4}&\frac{139}{12}\\\hdashline\frac{140}{3}&1&\frac{1}{4}&\frac{140}{12}\\\hline\end{array}"